OpenStudy (doc.brown):

How many plants can a gardener plant in a 25m^2 garden, when the plants must have a spacing of 0.2m?

OpenStudy (fibonaccichick666):

so, draw a picture to start

OpenStudy (doc.brown):



OpenStudy (doc.brown):

Is it 25m^2 = 5m * 5m 5m/0.2m = 25 plants Last plant doesn't count because you can't plant on the last edge. So 24*24 = 576?

OpenStudy (fibonaccichick666):

that sounds about right, the only thing I am concerned about is 1, the shape of the plants and garden, 2 in your pic, you should draw a square not a circle

OpenStudy (fibonaccichick666):

my other quandary is if the plants need .2 from the edge of the patch as well

OpenStudy (fibonaccichick666):

but, the last plant doesn't count, I disagree

OpenStudy (doc.brown):

Yes, they need a 0.2 space from anything, edge or plant

OpenStudy (fibonaccichick666):

ok, so, that means we don't actually have(assuming the garden is a square) a 5m side. we have to remove that edge distance which

OpenStudy (fibonaccichick666):

note, that if you would draw in all of the squares, the plants would fall on the end pts of the square,

OpenStudy (anonymous):